# Characterizations of zero singular ideal in étale groupoid C*-algebras via compressible maps

### v2

@@ 1. Introduction
Let $G$ be an étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)=\{0\}$ if and only if for every $x\in G^{0}$ and finite set $F\subseteq G^{x}_{x}$, the vectors $\delta_{fX\cap F}:=\sum_{h\in fX\cap F}\delta_{h}$ for $f\in F$ and $X\in\mathcal{X}(x)$ linearly span $\mathbb{C}[F]$.
The proof (of the contrapositive) of the “if” direction provides a new way to construct elements in the singular ideal. The spanning condition in Theorem F is also equivalent to the set of linear equations
@@
having no non-zero solution in $\mathbb{C}[G^{x}_{x}]$ (see Lemma 5.3). Following a similar proof to Corollary 5.20 and applying [5, Theorem 4.2], it is easy to see the Steinberg algebra of a ring $R$ and ample étale groupoid $G$ (see [28]) has zero singular ideal if and only if the above equations have no non-zero solution in $\mathbb{Z}_{t}[G^{x}_{x}]$ for every $x\in G^{0}$ and order $t\in\mathbb{N}\cup\{0\}$ of some non-zero element in $R$, where $\mathbb{Z}_{t}=\mathbb{Z}/t\mathbb{Z}$ for $t>0$ and $\mathbb{Z}_{0}=\mathbb{Q}$.
@@
We show these two questions always have the same answer.
@@
Theorem I (5.22 and 5.3).
Let $G$ be an étale groupoid such that $(G^{x}_{x},\mathcal{X}(x))\in\mathcal{I}$ for every $x\in G^{0}$. Then, $J=\{0\}$ if and only if for every $x\in G^{0}$, the set of linear equations
@@
has no non-zero solution in $\mathbb{C}[G^{x}_{x}]$.
@@ 5.3. Characterization of vanishing of the singular ideal
Let’s identify the isotropy fibres of $J\cap\mathscr{C}_{c}(G)$ using an explicit construction. We will use this later to characterize the largest class of étale groupoid C*-alebras (that can be defined in a certain way) for which the property $J\neq\{0\}\implies J\cap\mathscr{C}_{c}(G)\neq\{0\}$in terms of a property of the isotropy group C*-algebras. To see that it is the largest class, we will need the construction in Section 6, however.
@@
Let $G$ be an étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)=\{0\}$ if and only if for every $x\in D_{0}$ and finite set $F\subseteq G^{x}_{x}$, the vectors $\delta_{fX\cap F}:=\sum_{h\in fH\cap F}\delta_{h}$ for $f\in F$ and $X\in\pi^{-1}_{ess}(x)$ linearly span $\mathbb{C}[F]$.
Proof.
By Lemma 5.3 and Proposition 5.19, for $x\in D_{0}$, we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))\neq 0$ if and only if there is a finite set $F\subseteq G^{x}_{x}$ such that $\text{span}\{\delta_{fX\cap F}:f\in F,X\in\mathcal{X}\}\neq\mathbb{C}[F]$. Now, apply Lemma 5.8 to some non-zero element positive element in $J\cap\mathscr{C}_{c}(G)$ to show $J\cap\mathscr{C}_{c}(G)\neq 0$ if and only there is $x\in D_{0}$ such that $\eta_{x}(J\cap\mathscr{C}_{c}(G))\neq 0$, which proves the corollary. ∎
@@
Let $G$ be an étale groupoid.
Suppose $(G^{x}_{x},\pi^{-1}_{ess}(x))\in\mathcal{I}$ for every $x\in D_{0}$. Then, $J=\{0\}$ if and only if for every $x\in D_{0}$ and finite set $F\subseteq G^{x}_{x}$, the vectors $\delta_{fH\cap F}$ for $f\in F$ and $H\in\pi_{ess}^{-1}(x)$ linearly span $\mathbb{C}[F]$.
@@
Since the quotient map $C^{*}_{r(x)}(G^{x}_{x})\to C^{*}_{r\hat{\mathcal{X}}}(G^{x}_{x})$, where $\mathcal{X}=\pi^{-1}_{ess}(x)$ surjects $\text{ker}(q_{x})\cap\mathbb{C}[G^{x}_{x}]$ bijectively onto $J_{G^{x}_{x},\pi^{-1}_{ess}(x)}\cap\mathbb{C}[G^{x}_{x}]\neq\{0\}$ (trivially), it follows from Proposition 5.19 that $\eta_{x}(J\cap\mathscr{C}_{c}(G))\neq\{0\}$. Hence, $J\cap\mathscr{C}_{c}(G)\neq\{0\}$, proving the “if” direction.

### v3

@@ 1. Introduction
Let $G$ be an étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)=\{0\}$ if and only if for every $x\in G^{0}$ the set of linear equations
@@
has no non-zero finitely supported integer solution $(a_{h})_{h\in G^{x}_{x}}$.
For clarity, finitely supported means the integers $a_{h}=0$ for all but finitely many $h\in G^{x}_{x}$. We get several other characterizations of $J\cap\mathscr{C}_{c}(G)=\{0\}$ by applying Lemma 5.3 to Theorem 5.20, but we do not present these. The proof (of the contrapositive) of the “if” direction provides a new way to construct elements in the singular ideal.
Following a similar proof to Corollary 5.21 and applying [5, Theorem 4.2], it is easy to see the Steinberg algebra of a ring $R$ and ample étale groupoid $G$ (see [28]) has zero singular ideal if and only if the above equations in Theorem F have no non-zero solution in the group ring $\mathbb{Z}_{t}[G^{x}_{x}]$ for every $x\in G^{0}$ and order $t\in\mathbb{N}\cup\{0\}$ of some non-zero element in $R$, where $\mathbb{Z}_{t}=\mathbb{Z}/t\mathbb{Z}$ for $t>0$ and $\mathbb{Z}_{0}=\mathbb{Z}$.
@@
See Lemma 5.3 for different characterizations of this question. We show the above two questions always have the same answer.
@@
Theorem I (5.23 and 5.24).
Let $G$ be an étale groupoid such that $(G^{x}_{x},\mathcal{X}(x))\in\mathcal{I}$ for every $x\in G^{0}$. Then, $J=\{0\}$ if and only if $J\cap\mathscr{C}_{c}(G)=\{0\}$, if and only if for every $x\in G^{0}$, the set of linear equations
@@
has no non-zero finitely supported integer solution $(a_{h})_{h\in G^{x}_{x}}$.
@@
A countable abelian group $\Gamma$ satisfies Property $AI$ if and only if for every prime $p$, there is at most one element $g\in\Gamma$ with cyclic order $p$.
@@ 5.1. Coset groupoids
    There is a non-zero element $a\in\mathbb{Z}[\Gamma]$ such that $\sum_{x\in X}a(\gamma x)=0$ for all $\gamma\in\Gamma$ and $X\in\mathcal{X}$.
  4. (4)
@@
Obviously $(3)$ implies $(2)$, so let’s show the converse. Let $a\in\mathbb{C}[\Gamma]$ be as in $(2)$. Set $F=\{\gamma\in\Gamma:a(\gamma)\neq 0\}$, which is finite. Let $K$ be the collection of non-empty sets of the form $\gamma X\cap F$ for some $\gamma\in\Gamma$ and $X\in\mathcal{X}$. Then, each $\gamma X\cap F\in K$ can be written as $\gamma X\cap F=fX\cap F$ for some $f\in F$. Therefore, $K$ is finite. Let $A:\mathbb{C}^{F}\to\mathbb{C}^{K}$ be the matrix specified by $A_{\gamma,fX\cap F}=1$ if $\gamma\in fX\cap F$, and zero otherwise. Then, $a^{\prime}=(a(g))_{g\in F}\in\mathbb{C}^{F}$ satisfies $A(a^{\prime})=0$. Using Gaussian elimination, it is easy to see an integer valued matrix has a non-zero complex solution to $A(x)=0$ if and only if it has a non-zero rational solution $x=b^{\prime}$. By clearing denominators we may assume $b^{\prime}\in\mathbb{Z}^{F}$. Setting $b=\sum_{g\in F}b^{\prime}(g)\delta_{g}$, it is easy to see $b$ satisfies $(3)$.
@@ 5.3. Characterization of vanishing of the singular ideal
Let’s identify the isotropy fibres of $J\cap\mathscr{C}_{c}(G)$ using an explicit construction. We will use this later to characterize the largest class of étale groupoid C*-alebras (that can be defined in a certain way) for which the property $J\neq\{0\}\implies J\cap\mathscr{C}_{c}(G)\neq\{0\}$ holds in terms of a property of the isotropy group C*-algebras. To see that it is the largest class, we will need the construction in Section 6, however.
@@
Theorem 5.20.
Let $G$ be an étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)=\{0\}$ if and only if $J_{G^{x}_{x},\pi^{-1}_{ess}(x)}\cap\mathbb{C}[G^{x}_{x}]=0$ for every $x\in D_{0}$.
Proof.
By Proposition 5.19, we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))=J_{G^{x}_{x},\pi^{-1}_{ess}(x)}\cap\mathbb{C}[G^{x}_{x}]$. Now, apply Lemma 5.8 to some non-zero positive element in $J\cap\mathscr{C}_{c}(G)$ to show $J\cap\mathscr{C}_{c}(G)\neq 0$ if and only there is $x\in D_{0}$ such that $J_{G^{x}_{x},\pi^{-1}_{ess}(x)}\cap\mathbb{C}[G^{x}_{x}]=\eta_{x}(J\cap\mathscr{C}_{c}(G))\neq 0$, which proves the theorem. ∎
Now, we can apply our characterization of when $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]=\{0\}$ in Lemma 5.3 $(3)$ to immediately obtain the following corollary of Theorem 5.20.
@@
Let $G$ be an étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)=\{0\}$ if and only if for every $x\in D_{0}$, there is no non-zero $a\in\mathbb{Z}[G^{x}_{x}]$ such that
$\sum_{h\in gX}a(h)=0,\text{ for all }g\in G^{x}_{x},\text{ }X\in\pi^{-1}_{ess}(x).$
@@
Let $G$ be an étale groupoid, and assume $(G^{x}_{x},\pi^{-1}_{ess}(x))\in\mathcal{I}$ for every $x\in D_{0}$. Then, $J=\{0\}$ if and only if $J\cap\mathscr{C}_{c}(G)=\{0\}$.
@@
Since the quotient map $C^{*}_{r(x)}(G^{x}_{x})\to C^{*}_{r\hat{\mathcal{X}}}(G^{x}_{x})$, where $\mathcal{X}=\pi^{-1}_{ess}(x)$, surjects $\text{ker}(q_{x})\cap\mathbb{C}[G^{x}_{x}]$ bijectively onto $J_{G^{x}_{x},\pi^{-1}_{ess}(x)}\cap\mathbb{C}[G^{x}_{x}]\neq\{0\}$ (trivially), it follows from Proposition 5.19 that $\eta_{x}(J\cap\mathscr{C}_{c}(G))\neq\{0\}$. Hence, $J\cap\mathscr{C}_{c}(G)\neq\{0\}$, proving the “if” direction.
@@
We will verify the hypothesis for the above theorem for a large class of groupoids in Section 7. Theorem 5.23 combined with Corollary 5.21 yields the following.
Theorem 5.24.
Let $G$ be an étale groupoid and assume $(G^{x}_{x},\pi^{-1}_{ess}(x))\in\mathcal{I}$ for every $x\in D_{0}$. Then, $J=\{0\}$ if and only if for every $x\in D_{0}$, there is no non-zero $a\in\mathbb{Z}[G^{x}_{x}]$ such that
$\sum_{h\in gX}a(h)=0,\text{ for all }g\in G^{x}_{x},\text{ }X\in\pi^{-1}_{ess}(x).$